# On local Galois deformation rings: generalised reductive groups

### v1

@@ 15. Irreducible components and normality
The action of $G$ by conjugation on $\operatorname{Hom}_{\mathcal{O}\text{-}\mathrm{GrpSch}}(G^{0}/G^{\prime},\mathbb{G}_{m})=M$ induces an action of $\Delta$ on $M$. We assume that there is a $\Delta$-invariant decomposition
@@ 15.1. Dimension
Proposition 15.1 together with [52, Theorem 2 (ii)] imply that
@@
$0\rightarrow Z_{1}\cap Z_{2}G^{\prime}\rightarrow Z_{1}\rightarrow H_{1}^{0}\rightarrow 0$
@@
It follows from Corollary 15.4 that $Y^{\psi_{1}}_{\overline{\rho}^{\mathrm{ss}}}$ is the preimage of closed points in $X^{\mathrm{gen},\psi_{1}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$. Given this the proof is the same as the proof of [8, Lemma 3.21]. ∎
@@ 15.2. Generalised tori
is a finite abelian group of order $p^{m}$. We assume that $L$ contains all the $p^{m}$-th roots of unity. We further assume that $L$ is large enough so that $X^{\mathrm{gen}}_{H_{2},\overline{\psi}_{2}}(\mathcal{O})$ is non-empty. This is possible by [43, Theorem 9.3]. Under these assumptions we prove in [43, Theorem 9.3] that there is an isomorphism of complete local noetherian $\mathcal{O}$-algebras:
@@
where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{2}-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}(M_{2})_{\Delta}$. The Artin map of local class field theory $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ identifies $\mu$ with the torsion subgroup of $(\Gamma_{E}^{\mathrm{ab},p}\otimes M_{2})^{\Delta}$. We show in [43, Corollary 8.8] that
@@
Let $Z$ be a scheme over $\overline{X}^{\mathrm{ps}}_{H,\overline{\psi}^{\mathrm{ss}}_{2}}$ Then the underlying reduced subschemes of $Z$ and $Z^{\chi}$ coincide. In particular, $\dim Z=\dim Z^{\chi}$.
@@ 15.3. Serre’s criterion for normality
Since $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\rightarrow X^{\mathrm{gen}}_{\overline{G},\overline{\rho}^{\mathrm{ss}}}$ is finite by 15.1, the map $Z^{\operatorname{spcl}}_{G}\rightarrow Z^{\operatorname{spcl}}_{\overline{G}}$ is finite, and hence $\dim Z^{\operatorname{spcl}}_{G}\leq\dim Z^{\operatorname{spcl}}_{\overline{G}}$. Since $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}=\dim X^{\mathrm{gen}}_{\overline{G},\overline{\rho}^{\mathrm{ss}}}$ by 15.14 and 13.28 the assertion follows from Theorem 14.6 applied to $\overline{G}$. The argument for the special and generic fibres is the same. ∎
@@
It follows from [8, Lemma 2.1] that their connected components are $Z_{2}G^{\prime}$-invariant. If there was more than one then [52, Theorem 3 (iii)] implies that their images would disconnect $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$ (resp. $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$). If the orders of $(G/G^{0})(\overline{k})$ and $(Z_{1}\cap Z_{2}G^{\prime})(\overline{k})$ are coprime then this is not possible as the GIT-quotients are spectra of local rings in both cases by 15.4. ∎
@@
Let $A$ be either $A^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ or $A^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}/\varpi$. Since $A$ is of finite type over a complete local noetherian ring, it is excellent by [54, Tag 07QW]. Let $\mathfrak{p}$ be a prime of $A$. If $A$ is regular in codimension $i$ then $A_{\mathfrak{p}}$ is regular in codimension $i$, and since $A_{\mathfrak{p}}$ is again excellent, its completion $\hat{A}_{\mathfrak{p}}$ is also regular in codimension $i$ by [41, Theorem 23.9 (ii)]. (Since $A_{\mathfrak{p}}$ is excellent, it is a G-ring, and this implies that the assumptions made in the reference are satisfied. This is used implicitly in the proof of [41, Theorem 32.2 (i)].) By repeating the same argument with $A[T]$ we deduce that $\hat{A}_{\mathfrak{p}}\llbracket T\rrbracket$ is regular in codimension $i$. Lemma 15.17 and Lemma 5.17 imply that $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ and $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ are regular in codimension $c(\overline{\rho}^{\mathrm{ss}})$. Since $c(\overline{\rho}^{\mathrm{ss}})\geq 1$, Corollary 15.12 imply that Serre’s criterion for normality holds and we deduce that $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ and $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}/\varpi$ are normal domains. ∎
@@
Assume that either $[F:\mathbb{Q}_{p}]\geq 3$ or $[F:\mathbb{Q}_{p}]=2$ and there is no $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroup $L$ of $G$ such that $\dim G-\dim L=2$.
@@ 15.4. Irreducible components
Since $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is a subring of $A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ it is $\mathcal{O}$-torsion free (and hence $\mathcal{O}$-flat) and reduced. Hence, the irreducible components of $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ coincide. Since $A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is normal by Theorem 15.19 and $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]=(A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p])^{G^{0}}$, [13, Proposition 6.4.1] implies that $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is normal. Since $G^{0}$ is connected the irreducible components of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ are $G^{0}$-invariant and irreducible components of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ correspond to the GIT-quotients of irreducible components of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$. ∎
@@ 15.6. Labelling irreducible components
We show in [43, Section 7.4] that there is a canonical action of $\mathrm{X}(\mu)$ on the set of irreducible components of $X^{\mathrm{ps}}_{H_{2}}=\operatorname{Spec}R^{\mathrm{ps}}_{H_{2},\overline{\psi}^{\mathrm{ss}}_{2}}$, and this action is faithful and transitive. So to fix an $\mathrm{X}(\mu)$-equivariant bijection between $\mathrm{X}(\mu)$ and the set of irreducible components it is enough to specify to which component the identity in $\mathrm{X}(\mu)$ corresponds to. We don’t know how to do this canonically in general, except in one case, which we will now discuss.
@@
We will now give a different description of the bijection in the semi-direct product case. Let $X$ be an irreducible component of $X^{\mathrm{ps}}_{H_{2}}$. It follows from (93) and (95) that every point $x\in X^{\mathrm{ps}}_{H_{2}}(\overline{\mathbb{Q}}_{p})$ lies on a unique irreducible component and every irreducible component contains a ¯$\overline{\mathbb{Q}}_{p}$-valued point. We choose $x\in X(\overline{\mathbb{Q}}_{p})$ and let $\psi_{2}:\Gamma_{F}\rightarrow H_{2}(\overline{\mathbb{Q}}_{p})$ be a continuous representation such that its $H_{2}$-pseudocharacter corresponds to $x$. The representation $\psi_{2}$ is uniquely determined by $x$ up to $H_{2}^{0}(\overline{\mathbb{Q}}_{p})$-conjugation. Using that $H_{2}$ is a semi-direct product, we may write $\psi_{2}(\gamma)=(\Phi(\gamma),\gamma|_{E})$ for all $\gamma\in\Gamma_{F}$. Since $\psi_{2}$ is a continuous representation one may check that $\Phi:\Gamma_{F}\rightarrow H_{2}^{0}(\overline{\mathbb{Q}}_{p})$ defines a continuous $1$-cocycle and the $H_{2}^{0}(\overline{\mathbb{Q}}_{p})$-conjugacy class of $\psi_{2}$ defines a cohomology class in $H^{1}(\Gamma_{F},H^{0}(\overline{\mathbb{Q}}_{p}))$. We show in [43] using arguments that go back to the work of Langlands [37] on his correspondence for tori that
@@
The Artin map induces an isomorphism between $\mu:=(\mu_{p^{\infty}}(E)\otimes M_{2})^{\Delta}$ and the $p$-power torsion subgroup of $(\Gamma_{F}^{\mathrm{ab},p}\otimes M_{2})^{\Delta}$. Let $\chi:\mu\rightarrow\overline{\mathbb{Q}}_{p}^{\times}$ be the restriction of $\tilde{\chi}$ to $\mu$. Then $\chi$ takes values in $\mathcal{O}^{\times}\subset\overline{\mathbb{Q}}_{p}^{\times}$ and we show in [43, Section 7.4] that the irreducible component $X$ is uniquely determined by $\chi$.
@@ 16. Deformations into $L$-groups and $C$-groups
Let $H$ be a connected reductive group defined over $F$ and let $E$ be a finite Galois extension of $F$ such that $H_{E}$ is split. Let $T$ be a maximal split torus in $H_{E}$ and let $B$ be a Borel subgroup of $H_{E}$ containing $T$. To the triple $(H_{E},B,T)$ one may attach a based root datum $(X^{\ast}(T),\Phi,\Delta,X_{\ast}(T),\Phi^{\vee},\Delta^{\vee})$, where the symbols denote characters of $T$, the set of all roots, the set of positive roots corresponding to $B$, cocharacters of $T$, the set of coroots, and the set of positive coroots corresponding to $B$, respectively, see [20, Section 1.5]. The dual group is a triple $(\widehat{H},\widehat{B},\widehat{T}$), where $\widehat{H}$ is a connected split reductive group scheme defined over $\mathbb{Z}$, $\widehat{B}$ is a Borel subgroup scheme of $\widehat{H}$ containing a maximal split torus $\widehat{T}$, such that the based root datum of $(\widehat{H},\widehat{B},\widehat{T})$ is isomorphic to $(X_{\ast}(T),\Phi^{\vee},\Delta^{\vee},X^{\ast}(T),\Phi,\Delta)$. Out of the action of $\operatorname{Gal}(E/F)$ on $H_{E}$ one may construct a group of homomorphism of $\operatorname{Gal}(E/F)$ to the automorphism group of the based root datum. After fixing additional datum, called a pinning, one rigidifies the situation to obtain a group homomorphism $\operatorname{Gal}(E/F)\rightarrow\operatorname{Aut}(\widehat{H})$. We refer the reader to [10, Section I.1], or [23, Section 2] for further details. The $L$-group is defined as
@@
The base change of $Z(H)^{0}$ to $E$ is equal to $Z(H_{E})^{0}$. It follows from [11, Proposition 14.2] that
@@
Thus the canonical identification $X_{\ast}(T)=X^{\ast}(\widehat{T})$ identifies $X_{\ast}(Z(H_{E})^{0})$ with $M$. Since this identification is $\operatorname{Gal}(E/F)$-equivariant we obtain
@@
Let us make the bijection more explicit. If $X$ is a an irreducible component of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ then we may pick a geometric point $y\in X(\overline{\mathbb{Q}}_{p})$. Then $y$ corresponds to a continuous representation $\rho:\Gamma_{F}\rightarrow{}^{L}H(\overline{\mathbb{Q}}_{p})$. By quotienting out $\widehat{H}^{\prime}(\overline{\mathbb{Q}}_{p})$ we obtain a continuous admissible representation
@@
The group ${}^{L}H/\widehat{H}^{\prime}$ is the $L$-group for $Z(H)^{0}(F)$. The $\widehat{H}(\overline{\mathbb{Q}}_{p})$-conjugacy of $\psi$ corresponds to a continuous character $\tilde{\chi}:Z(H)^{0}(F)\rightarrow\overline{\mathbb{Q}}_{p}$ under the local Langlands correspondence for tori, [37], [6]. By restricting $\tilde{\chi}$ to $Z(H)^{0}(F)_{p^{\infty}}$ we obtain the required character $\chi$.
@@
The $C$-group ${}^{C}H$ of $H$ is a variant of an $L$-group. It has been defined in [14, Definition 5.3.2]; a nice exposition is given in [61, Section 1.1]. For us only the following properties are important: there is an exact sequence
[…]
@@
Let $D^{C}_{\overline{\rho}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor such that $D^{C}_{\overline{\rho}}(A)$ is the set of continuous representations $\rho_{A}:\Gamma_{F}\rightarrow{}^{C}H(A)$, such that $\rho_{A}\equiv\overline{\rho}\pmod{\mathfrak{m}_{A}}$ and $d\circ\rho_{A}=\chi_{\mathrm{cyc}}\otimes_{\mathcal{O}}A$. The definition of this deformation problem is motivated by the first two bullet points in [14, Conjecture 5.3.4], which describes Galois representations attached to $C$-algebraic automorphic forms. The functor $D^{C}_{\overline{\rho}}$ is representable by a complete local noetherian $\mathcal{O}$-algebra $R^{\square,C}_{\overline{\rho}}$ with residue field $k$.
@@
We then proceed as in the proof of Theorem 16.4. The group ${}^{L}H/\widehat{H}^{\prime}$ is the $L$-group for $Z(H)^{0}(F)$. The $\widehat{H}(\overline{\mathbb{Q}}_{p})$-conjugacy of $\psi_{2}$ corresponds to a continuous character $\tilde{\chi}:Z(H)^{0}(F)\rightarrow\overline{\mathbb{Q}}_{p}$ under the local Langlands correspondence for tori, [37], [6]. By restricting $\tilde{\chi}$ to $Z(H)^{0}(F)_{p^{\infty}}$ we obtain the required character $\chi$. ∎

### v2

@@ 15. Irreducible components and normality
The action of $G$ by conjugation on $\operatorname{Hom}_{\mathcal{O}\text{-}\mathrm{GrpSch}}(G^{0}/G^{\prime},\mathbb{G}_{m})=M$ induces an action of $\Delta$ on $M$. We assume that we are given a $\Delta$-invariant decomposition
@@ 15.1. Dimension
Proposition 15.1 together with [57, Theorem 2 (ii)] imply that
@@
$0\rightarrow Z_{1}\cap Z_{2}G^{\prime}\rightarrow Z_{1}\rightarrow H_{1}^{0}\rightarrow 0.$
@@
It follows from Corollary 15.4 that $Y^{\psi_{1}}_{\overline{\rho}^{\mathrm{ss}}}$ is the preimage of closed points in $X^{\mathrm{gen},\psi_{1}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$. Given this the proof is the same as the proof of [6, Lemma 3.21]. ∎
@@ 15.2. Generalised tori
is a finite abelian group of order $p^{m}$. We assume that $L$ contains all the $p^{m}$-th roots of unity. We further assume that $L$ is large enough so that $X^{\mathrm{gen}}_{H_{2},\overline{\psi}_{2}}(\mathcal{O})$ is non-empty. This is possible by [48, Theorem 9.3]. Under these assumptions we prove in [48, Theorem 9.3] that there is an isomorphism of complete local noetherian $\mathcal{O}$-algebras:
@@
where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{2}-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}(M_{2})_{\Delta}$. The Artin map of local class field theory $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ identifies $\mu$ with the torsion subgroup of $(\Gamma_{E}^{\mathrm{ab},p}\otimes M_{2})^{\Delta}$. We show in [48, Corollary 8.8] that
@@
Let $Z$ be an affine scheme over $\overline{X}^{\mathrm{ps}}_{H,\overline{\psi}^{\mathrm{ss}}_{2}}$ Then the underlying reduced subschemes of $Z$ and $Z^{\chi}$ coincide. In particular, $\dim Z=\dim Z^{\chi}$.
@@ 15.3. Serre’s criterion for normality
Since $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\rightarrow X^{\mathrm{gen}}_{\overline{G},\overline{\rho}^{\mathrm{ss}}}$ is finite by 15.1, the map $Z^{\operatorname{spcl}}_{G}\rightarrow Z^{\operatorname{spcl}}_{\overline{G}}$ is finite, and hence $\dim Z^{\operatorname{spcl}}_{G}\leq\dim Z^{\operatorname{spcl}}_{\overline{G}}$. Since $\dim X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}=\dim X^{\mathrm{gen}}_{\overline{G},\overline{\rho}^{\mathrm{ss}}}$ by 15.14 and 13.28 the assertion follows from Theorem 14.6 applied to $\overline{G}$. The argument for the special and generic fibres is the same. ∎
@@
It follows from [6, Lemma 2.1] that their connected components are $Z_{2}G^{\prime}$-invariant. If there was more than one then [57, Theorem 3 (iii)] implies that their images would disconnect $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$ (resp. $X^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash Z_{2}G^{\prime}$). If the orders of $(G/G^{0})(\overline{k})$ and $(Z_{1}\cap Z_{2}G^{\prime})(\overline{k})$ are coprime then this is not possible as the GIT quotients are spectra of local rings in both cases by 15.4. ∎
@@
Let $A$ be either $A^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ or $A^{\mathrm{gen},\psi_{1},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}/\varpi$. Since $A$ is of finite type over a complete local noetherian ring, it is excellent by [59, Tag 07QW]. Let $\mathfrak{p}$ be a prime of $A$. If $A$ is regular in codimension $i$ then $A_{\mathfrak{p}}$ is regular in codimension $i$, and since $A_{\mathfrak{p}}$ is again excellent, its completion $\hat{A}_{\mathfrak{p}}$ is also regular in codimension $i$ by [43, Theorem 23.9 (ii)]. (Since $A_{\mathfrak{p}}$ is excellent, it is a G-ring, and this implies that the assumptions made in the reference are satisfied. This is used implicitly in the proof of [43, Theorem 32.2 (i)].) By repeating the same argument with $A[T]$ we deduce that $\hat{A}_{\mathfrak{p}}\llbracket T\rrbracket$ is regular in codimension $i$. Lemma 15.17 and Lemma 5.17 imply that $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ and $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ are regular in codimension $c(\overline{\rho}^{\mathrm{ss}})$. Since $c(\overline{\rho}^{\mathrm{ss}})\geq 1$, Corollary 15.12 imply that Serre’s criterion for normality holds and we deduce that $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}$ and $R^{\square,\psi_{1},\chi}_{G,\rho_{x}}/\varpi$ are normal domains. ∎
@@
Assume that either $F:\mathbb{Q}_{p}]\geq 3$ or $F:\mathbb{Q}_{p}]=2$ and there is no $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroup $L$ of $G$ such that $\dim G-\dim L=2$.
@@ 15.4. Irreducible components
Since $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is a subring of $A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ it is $\mathcal{O}$-torsion free (and hence $\mathcal{O}$-flat) and reduced. Hence, the irreducible components of $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ coincide. Since $A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is normal by Theorem 15.19 and $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]=(A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p])^{G^{0}}$, [12, Proposition 6.4.1] implies that $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is normal. Since $G^{0}$ is connected the irreducible components of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ are $G^{0}$-invariant and since $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is normal its irreducible and connected components coincide. Thus irreducible components of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ correspond to the GIT quotients of irreducible components of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$. ∎
@@ 15.6. Labelling irreducible components
We show in [48, Section 7.4] that there is a canonical action of $\mathrm{X}(\mu)$ on the set of irreducible components of $X^{\mathrm{ps}}_{H_{2}}=\operatorname{Spec}R^{\mathrm{ps}}_{H_{2},\overline{\psi}^{\mathrm{ss}}_{2}}$, and this action is faithful and transitive. So to fix an $\mathrm{X}(\mu)$-equivariant bijection between $\mathrm{X}(\mu)$ and the set of irreducible components it is enough to specify to which component the identity in $\mathrm{X}(\mu)$ corresponds to. We don’t know how to do this canonically in general, except in one case, which we will now discuss.
@@
We will now give a different description of the bijection in the semi-direct product case. Let $X$ be an irreducible component of $X^{\mathrm{ps}}_{H_{2}}$. It follows from (93) and (95) that every point $x\in X^{\mathrm{ps}}_{H_{2}}(\overline{\mathbb{Q}}_{p})$ lies on a unique irreducible component and every irreducible component contains a $\overline{\mathbb{Q}}_{p}$-valued point. We choose $x\in X(\overline{\mathbb{Q}}_{p})$ and let $\psi_{2}:\Gamma_{F}\rightarrow H_{2}(\overline{\mathbb{Q}}_{p})$ be a continuous representation such that its $H_{2}$-pseudocharacter corresponds to $x$. The representation $\psi_{2}$ is uniquely determined by $x$ up to $H_{2}^{0}(\overline{\mathbb{Q}}_{p})$-conjugation. Using that $H_{2}$ is a semi-direct product, we may write $\psi_{2}(\gamma)=(\Phi(\gamma),\gamma|_{E})$ for all $\gamma\in\Gamma_{F}$. Since $\psi_{2}$ is a continuous representation one may check that $\Phi:\Gamma_{F}\rightarrow H_{2}^{0}(\overline{\mathbb{Q}}_{p})$ defines a continuous $1$-cocycle and the $H_{2}^{0}(\overline{\mathbb{Q}}_{p})$-conjugacy class of $\psi_{2}$ defines a cohomology class in $H^{1}(\Gamma_{F},H^{0}(\overline{\mathbb{Q}}_{p}))$. We show in [48] using arguments that go back to the work of Langlands [38] on his correspondence for tori that
@@
The Artin map induces an isomorphism between $\mu:=(\mu_{p^{\infty}}(E)\otimes M_{2})^{\Delta}$ and the $p$-power torsion subgroup of $(\Gamma_{F}^{\mathrm{ab},p}\otimes M_{2})^{\Delta}$. Let $\chi:\mu\rightarrow\overline{\mathbb{Q}}_{p}^{\times}$ be the restriction of $\tilde{\chi}$ to $\mu$. Then $\chi$ takes values in $\mathcal{O}^{\times}\subset\overline{\mathbb{Q}}_{p}^{\times}$ and we show in [48, Section 7.4] that the irreducible component $X$ is uniquely determined by $\chi$.
@@ 16. Deformations into $L$-groups and $C$-groups
[…]
@@
The base change of $Z(H)^{0}$ to $E$ is equal to $Z(H_{E})^{0}$. It follows from [10, Proposition 14.2] that
@@
[…]
@@
Let us make the bijection more explicit. If $X$ is an irreducible component of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ then we may pick a geometric point $y\in X(\overline{\mathbb{Q}}_{p})$. Then $y$ corresponds to a continuous representation $\rho:\Gamma_{F}\rightarrow{}^{L}H(\overline{\mathbb{Q}}_{p})$. By quotienting out $\widehat{H}^{\prime}(\overline{\mathbb{Q}}_{p})$ we obtain a continuous admissible representation
@@
The group $^{}^{L}H/\widehat{H}^{\prime}$ is the $L$-group for $Z(H)^{0}(F)$. The $\widehat{H}(\overline{\mathbb{Q}}_{p})$-conjugacy of $\psi$ corresponds to a continuous character $\tilde{\chi}:Z(H)^{0}(F)\rightarrow\overline{\mathbb{Q}}_{p}$ under the local Langlands correspondence for tori, [38], [4]. By restricting $\tilde{\chi}$ to $Z(H)^{0}(F)_{p^{\infty}}$ we obtain the required character $\chi$.
@@
The $C$-group $^{C}H$ of $H$ is a variant of an $L$-group. It has been defined in [13, Definition 5.3.2]; a nice exposition is given in [66, Section 1.1]. For us only the following properties are important: there is an exact sequence
(109) $0\rightarrow{}^{L}H\rightarrow{}^{C}H\overset{d}{\rightarrow}\mathbb{G}_{m}\rightarrow 0.$
@@
Let $D^{C}_{\overline{\rho}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor such that $D^{C}_{\overline{\rho}}(A)$ is the set of continuous representations $\rho_{A}:\Gamma_{F}\rightarrow{}^{C}H(A)$, such that $\rho_{A}\equiv\overline{\rho}\pmod{\mathfrak{m}_{A}}$ and $d\circ\rho_{A}=\chi_{\mathrm{cyc}}\otimes_{\mathcal{O}}A$. The definition of this deformation problem is motivated by the first two bullet points in [13, Conjecture 5.3.4], which describes Galois representations attached to $C$-algebraic automorphic forms. The functor $D^{C}_{\overline{\rho}}$ is representable by a complete local noetherian $\mathcal{O}$-algebra $R^{\square,C}_{\overline{\rho}}$ with residue field $k$.
@@
We then proceed as in the proof of Theorem 16.4. The group $^{}^{L}H/\widehat{H}^{\prime}$ is the $L$-group for $Z(H)^{0}(F)$. The $\widehat{H}(\overline{\mathbb{Q}}_{p})$-conjugacy of $\psi_{2}$ corresponds to a continuous character $\tilde{\chi}:Z(H)^{0}(F)\rightarrow\overline{\mathbb{Q}}_{p}$ under the local Langlands correspondence for tori, [38], [4]. By restricting $\tilde{\chi}$ to $Z(H)^{0}(F)_{p^{\infty}}$ we obtain the required character $\chi$. ∎